# CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?

### v1

@@ 3.3 Task C: Instruction Generation
The third task evaluates a model’s ability to generate natural-language crochet instructions directly from an image of a finished product. We construct CrochetBench-C, a subset of 6,009 examples, where each product image is paired with its full pattern. Unlike classification or selection, this task requires producing a multi-step sequence that follows domain-specific syntax, ordering, and stitch logic. Outputs are expected to resemble real-world crochet patterns, written line by line (e.g., “Rnd 1: ch 4, 6 sc in ring”). This setting challenges models to translate visual evidence into coherent procedural text that preserves both symbolic accuracy and structural consistency.
@@
We evaluate generation quality using standard text-generation metrics and a domain-specific structural measure. BLEU and ROUGE-L capture lexical and n-gram overlap with the reference instructions. ChrF, computed over character n-grams, provides a finer-grained signal of similarity and is particularly suited to stitch abbreviations. While these metrics measure surface-level overlap, they do not directly test whether the generated instructions are structurally consistent (e.g., balanced stitch counts, valid round progression). To address this limitation, we complement Task C with Task D, where compilation-based evaluation of DSL translations directly assesses executable consistency. Together, these tasks allow us to probe both linguistic fidelity and procedural correctness.
Table 5: Evaluation results on the Instruction Generation task. We report BLEU, ROUGE-L, and ChrF. Higher is better. Best results are bold; second-best are underlined.
Model Size BLEU ROUGE-L ChrF
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.0021 0.0926 9.32
Google Gemma 3 4B 0.0010 0.0329 5.17
Qwen2-VL 7B 0.0160 0.2084 15.76
DeepSeek-VL 7B 0.0133 0.1968 18.12
Closed Source GPT-4o 0.0333 0.2353 23.80
Gemini 2.5 Flash-Lite 0.0482 0.2583 30.20
Claude Sonnet 4 0.0331 0.2516 22.95
Results. Gemini achieves the strongest overall performance across all metrics, substantially outperforming other closed-source models. GPT-4o and Claude follow closely, while open-source models lag behind with significantly lower BLEU and ChrF scores. This gap highlights the difficulty of generating structurally faithful crochet instructions, which requires models to capture both visual details and domain-specific procedural logic.
@@ 3.4 Task D: Instruction-to-DSL Translation
We construct two subsets for Task D: CrochetBench-Dstep (119 items) and CrochetBench-Dproj (100 items). For CrochetBench-Dstep, we manually annotate the correct CrochetPARADE DSL for the previous context of each pattern (a prefix of NL (natural language) –DSL pairs). The model is then given this prefix along with the next natural-language instruction, and must predict the corresponding DSL line. Predictions are evaluated by checking compilation validity with the CrochetPARADE validator. For CrochetBench-Dproj, we adopt a few-shot prompting setup rather than full manual annotation. Annotators provide a single reference program in CrochetPARADE as context, which is paired with a new natural-language instruction and its product image. The model must then generate a complete DSL program, and outputs are assessed by whether they compile fully or partially using the validator.
Design Rationale. Crochet patterns admit many valid DSL realizations (e.g., alternative groupings or equivalent constructs), meaning that no single gold reference is canonical. A reference-based metric could unfairly penalize models that produce semantically correct but structurally different programs. By instead relying on the validator, CrochetBench shifts evaluation away from surface-level string matching toward functional executability—the central criterion for crochet synthesis. This design is consistent with other program synthesis benchmarks, such as semantic parsing and SQL generation, where execution accuracy is often preferred over exact string match.
3.4.1 Step-Level Translation
In the step-level setting, the model receives a prefix of NL–DSL pairs and must generate the DSL line corresponding to the next natural-language instruction. This setup reflects an incremental synthesis process in which correctness depends on maintaining stitch-level consistency across steps. Since crochet patterns are inherently stateful, earlier context is critical for resolving constructs such as increases, repeats, and turning chains. To capture progression through a pattern, we sample 52 early examples (steps 1–2), 34 mid examples (steps 3–4), and 33 late examples (steps 5–6). Step-level inputs are formatted as:
$\text{Prefix (NL--DSL pairs)}+\text{Next NL instruction}\;\;\rightarrow\;\;\text{Next DSL line}.$
This formulation allows models to generate locally plausible DSL steps, but global correctness ultimately requires stronger contextual reasoning.
@@ 3.4.1 Step-Level Translation
We evaluate models using Compilation Success Rate (CSR), defined as the proportion of generated DSL outputs that compile successfully with the CrochetPARADE validator. This validator-based metric is a key strength of CrochetBench: unlike reference-based string matching, CSR directly measures whether generated programs are executable, rewarding functional correctness rather than surface similarity. To better understand systematic errors, we also conduct fine-grained error analysis across four categories: (1) syntax structure errors, (2) stitch definition errors, (3) labeling and reference errors, and (4) structural or formatting issues. A detailed taxonomy with illustrative examples is provided in the Appendix.
CrochetBench further supports multimodal verification, as CrochetPARADE programs can be rendered into simulated product images. These renderings can be compared with ground-truth product images using pretrained vision–language models (e.g., CLIP), enabling evaluation of both structural validity and visual faithfulness. While current model performance leaves ample room for progress, the combination of validator-based execution checks and prospective image-grounded verification establishes CrochetBench as a uniquely rigorous testbed for structured multimodal reasoning.
Table 6: Verification results across models, grouped by category.
Model CSR (%) Undef. Br. MRef Other Tot. Err.
Open Source
Salesforce BLIP-2 Flan-T5 XL 4.2 29.8 48.2 20.2 1.8 114
Google Gemma 3 3.4 26.1 10.4 63.5 115
DeepSeek-VL 32.8 36.2 38.8 11.2 13.8 80
Qwen2-VL 35.3 42.9 39.0 1.3 16.8 77
Closed Source
GPT-4o 36.1 39.5 43.4 17.1 76
Gemini 2.5 Flash-Lite 41.2 51.4 21.4 5.7 21.5 70
Claude Sonnet 4 52.1 45.6 28.1 26.3 57
Results. Step-level translation is locally tractable: models often produce syntactically valid lines, but compilation failures reveal semantic underspecification. Claude achieves the highest valid rate (52.1%), while DeepSeek-VL and Qwen2-VL are the strongest among open-source models. Errors are dominated by syntax (brackets, references) and undefined stitches, underscoring the difficulty of maintaining state consistency even when local syntax is correct.
3.4.2 Project-Level Translation
In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program. This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design. Image grounding plays a crucial role in disambiguating constructs such as repeated motifs, symmetry, and termination conditions.
Evaluation.
We assess model outputs using two complementary metrics. The first is Compilation Success Rate (CSR), and the second is Partial Executable Rate (PER), which measures the average fraction of a program that compiles successfully before failure. While CSR captures all-or-nothing executability, PER provides a finer-grained view of structural alignment, offering credit to models that generate correct prefixes even if the full program does not compile. Error types are categorized using the same taxonomy as in the step-level evaluation.
Table 7: Project-level verification results. We report Compilation Success Rate (CSR) and Partial Executable Rate (PER). Best scores are bold, second-best are underlined.
Model Size CSR (%) PER (%)
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 1.0 0.00
Google Gemma 3 4B 1.6 5.29
DeepSeek-VL 7B 8.1 37.49
Qwen2-VL 7B 21.0 30.28
Closed Source GPT-4o 4.0 2.76
Gemini 2.5 Flash-Lite 4.0 5.67
Claude Sonnet 4 5.0 8.16
Table 8: Project-level error type distribution across models (percentage). Abbreviations: Undef. (Undefined), Br. (Brackets), Lbl. (Labels Missing), Non-adj. (Non-adjacent), MRef (Multi-Reference), Other (Other/Runtime).
Model Undef. Br. Lbl. Non-adj. MRef Other
Open Source
Salesforce BLIP-2 Flan-T5 XL 2.0 37.4 58.6 2.0
Google Gemma 3 39.3 14.8 11.5 8.2 14.8 11.5
DeepSeek-VL 36.3 45.1 4.4 11.0 3.3
Qwen2-VL 25.0 12.5 7.8 51.6 3.1
Closed Source
GPT-4o 61.5 7.3 16.7 4.2 4.2 6.2
Gemini 2.5 Flash-Lite 51.1 18.1 9.6 8.5 4.3 8.6
Claude Sonnet 4 46.3 14.7 21.1 7.4 5.3 5.3
Results. Project-level translation remains highly challenging. Qwen2-VL achieves the best valid rate (21.0%) and strong PER, surpassing all closed-source systems. DeepSeek-VL demonstrates robust partial executability despite a lower valid rate. By contrast, GPT-4o, Gemini, and Claude achieve lower scores, highlighting that even state-of-the-art VLMs struggle with long-range structural consistency in executable synthesis. Error analysis (Table 8) reveals that closed-source models often fail due to label and reference inconsistencies, while open-source models more frequently exhibit undefined stitches and syntax errors.
4 Discussion
The results across CrochetBench highlight both the promise and current limitations of Visual language models in bridging perception, procedural reasoning, and executable synthesis. By structuring tasks in a progressive ladder, we expose clear gradients of difficulty: while contemporary models demonstrate competence in low-level recognition and mid-level comprehension, their performance declines substantially when asked to generate or formalize instructions into compilable domain-specific programs. This pattern underscores both methodological bottlenecks in multimodal grounding and fundamental challenges in symbolic reasoning over long-horizon structures.
In Task A, stitch recognition is feasible for both open- and closed-source models, though precision–recall tradeoffs differ. Claude Sonnet 4 prioritizes precision for a higher F1, while DeepSeek-VL favors recall via richer visual encoders. This reflects divergent inductive biases: commercial models better regularize spurious predictions, while open-source models overpredict to capture subtle textures. Crucially, this is a multi-label, fine-grained texture classification task with symbolic implications, distinguishing it from generic object recognition. Baselines like BLIP-2, designed for captioning, underperform due to insufficient symbolic grounding (blip; clip). Task B amplifies these challenges. Closed-source models, especially GPT-4o, outperform open models by a wide margin, underscoring the difficulty of aligning visual cues with procedural semantics. Distractor instructions are intentionally plausible, demanding reasoning over local–global coherence, not just token overlap (GVE; VQA; miech2019howto100mlearningtextvideoembedding). Qwen2-VL’s competitive performance suggests that scale and pretraining diversity help, but commercial systems benefit from stronger instruction tuning and better vision–language alignment.
Task C exposes the steep drop in open-ended procedural generation. Gemini 2.5 leads across BLEU, ROUGE, and ChrF (papineni2002bleu; lin2004rouge; popovic2015chrf), reflecting fluency in structured text. Yet absolute scores are low, and outputs often fail to conform to crochet-specific syntax. Open models hallucinate frequently, lacking exposure to domain-aligned distributions. This highlights a broader issue: fluency in general language modeling does not imply competence in structured, domain-specific generation (miech2019howto100mlearningtextvideoembedding; alayrac2022flamingovisuallanguagemodel).
Task D proves most demanding. Even top closed-source models rarely exceed 6% project-level compilation success, while Qwen2-VL unexpectedly achieves 21%, suggesting better symbolic generalization under execution constraints. Syntax errors (e.g., undefined stitches, malformed brackets) dominate in open models, while closed models produce syntactically valid but semantically inconsistent programs. These complementary failure modes hint at differing generalization paths: symbolic robustness versus lexical fluency. Notably, Task D also tests 3D-aware procedural fidelity—models must translate visual or textual cues into symbolic programs that unfold into spatially coherent structures, not just grammatically valid sequences.
Taken together, these results demonstrate that success in linguistic generation does not translate directly to executable synthesis. Performance decays sharply when evaluation moves from surface-level fidelity (BLEU, ROUGE) to structural validity (compilation), reinforcing the importance of execution-grounded metrics for procedural tasks. The relative strength of Qwen2-VL at the project level further suggests that progress may come not from scaling alone, but from architectural or training adjustments that better capture long-range dependencies and stateful operations. More broadly, CrochetBench exposes a critical frontier for multimodal reasoning: models must not only ground text in vision but also internalize procedural invariants that guarantee functional correctness.
However, our analysis must be contextualized within certain constraints. The benchmarks reflect standardized stitch sets and normalized instructions, which, while ensuring comparability, simplify the variability encountered in real-world crochet practice (e.g., designer-specific shorthand, unconventional repeats). Moreover, evaluation metrics in Task C rely on string overlap, which may underestimate semantically correct but lexically divergent outputs. Even the compilation-based metric in Task D, though stronger, cannot assess visual fidelity unless paired with image-render verification. Finally, the sample size for project-level DSL translation remains limited, and absolute validity rates are low, constraining fine-grained statistical comparisons.
5 Future Work
A central avenue for future research is advancing models that translate free-form natural language crochet instructions into the formal CrochetPARADE DSL. As a domain-specific programming language, CrochetPARADE not only enables executable verification of multimodal outputs but also positions crochet as a program synthesis problem, where compilers map symbolic grammars into machine-executable instructions, extending ideas from prior DSL work in domains such as knitting (knitscript), graphics (ellis2018learninginfergraphicsprograms), and robotics (mu2024robocodexmultimodalcodegeneration). This perspective connects naturally to CAD/CAM integration: industrial crochet and warp-knitting machines (e.g., COMEZ, Jakob Müller) already rely on pipelines from graphical design interfaces, to intermediate graph-based representations, to low-level machine code. CrochetPARADE could serve as a standardized intermediate layer in this workflow, bridging human-facing authoring tools with machine-facing execution systems (text2cad).
Several technical directions arise from this framing. First, incorporating explicit state-tracking mechanisms—whether through memory-augmented architectures (graves2016hybrid) or symbolic scaffolds (nye2021workscratchpadsintermediatecomputation)—could mitigate long-range inconsistencies in DSL translation. Second, multimodal pretraining enriched with procedural and topological domains (e.g., assembly instructions (cao2023instruction), instructional videos (miech2019howto100mlearningtextvideoembedding)) may narrow the gap between natural-language fluency and executable synthesis. Third, hybrid evaluation pipelines that combine compilation checks with visual render comparisons could more holistically assess structural and perceptual fidelity, building on metrics used in programmatic 3D generation (jain2022zero; paschalidou2021neuralpartslearningexpressive). Finally, CrochetBench offers a testbed for exploring neuro-symbolic integration, where neural perception is paired with symbolic reasoning to reconcile fine-grained visual cues with globally coherent program execution.
In sum, CrochetBench and CrochetPARADE highlight a pronounced frontier at the intersection of multimodal learning, program synthesis, and digital fabrication. Progress here may one day close the loop from human intent to automated textile production, advancing both structured multimodal reasoning and computational craft.
Ethics Statement

### v2

@@ 3.3 Task C: Instruction Generation
Task C evaluates a model’s ability to generate natural-language crochet instructions from an image of a finished item. We construct CrochetBench-C, a subset of 6,009 examples in which each image is paired with the corresponding ground-truth textual instruction. In contrast to captioning or stylistic description, this task requires generating a sequence of domain-specific commands (e.g., “Rnd 1: ch 4, 6 sc in ring”), each of which encodes precise stitch operations, counts, and ordering. Because real crochet patterns may include tens of steps, hierarchical structure (rounds, rows, substeps), and long-range dependencies, this task assesses whether models can infer the underlying procedural logic implied by the final visual product. The generated text must maintain consistent stitch semantics, preserve temporal ordering, and follow established formatting conventions used by human crafters.
@@
We evaluate generation quality using BLEU, ROUGE-L, and ChrF (Papineni et al., 2002; Lin, 2004; Popović, 2015), which together capture complementary aspects of textual fidelity in procedural instructions. BLEU measures overlap of word-level n-grams and thus reflects local lexical accuracy in stitch tokens and command sequences. ROUGE-L evaluates the longest common subsequence between the generated and reference patterns, capturing larger-scale ordering and structural alignment across multi-step procedures. ChrF operates on character-level n-grams, which makes it effective for crochet patterns where stitch abbreviations (e.g., sc, sc2tog) often differ by only a few characters. Word-based metrics treat such tokens as entirely distinct, whereas character-level comparisons can capture partial matches and small but semantically important variations.
However, textual overlap metrics alone cannot reveal whether the generated instructions form a coherent or executable procedure. A model may generate instructions that appear fluent and pattern-like while still violating fundamental structural constraints, including inconsistent stitch counts, infeasible transitions, or unbalanced repeat constructions. To directly assess structural correctness and program-level understanding, we introduce Task D, which requires models to formalize correct natural-language instructions into a machine-checkable DSL representation.
@@ 3.4 Task D: Instruction-to-DSL Translation
Tasks A–C evaluate perception, retrieval, and natural-language generation, but they do not test whether a model can represent crochet procedures in a structured, machine-interpretable form. Crochet patterns are inherently programmatic: they contain loops, repeats, and counting logic that natural language expresses only implicitly, and that text-based metrics cannot reliably validate. Task D isolates this structural dimension by requiring models to translate correct natural-language instructions into an executable DSL, thereby revealing whether models grasp the underlying program-like structure of crochet. This capability is essential for true procedural reasoning, and we instantiate it using the CrochetPARADE DSL.
We construct two variants of Task D: CrochetBench-Dstep (119 items) for step-level formalization and CrochetBench-Dproj (100 items) for project-level program synthesis.
Step-Level Translation
The step-level task evaluates incremental translation from natural-language crochet instructions to DSL, where each instruction updates a shared stitch state. In this setting, the model is given a prefix of correct NL–DSL pairs representing the pattern translated so far and must generate the corresponding DSL line for the next instruction. Because crochet patterns are stateful, correct translation requires maintaining consistency with previous steps, including stitch counts, repeat structures, turning logic, and round-to-round topology.
This formulation tests whether models can map local textual cues into the structured symbolic operations of CrochetPARADE while preserving long-range procedural context. To capture variation in pattern progression, CrochetBench-Dstep groups examples by prefix length: 52 early-step examples use prefixes covering steps 1–2, 34 mid-step examples use prefixes covering steps 1–4, and 33 late-step examples use prefixes covering steps 1–6. Each example requires translating the next natural-language instruction given the corresponding NL–DSL history.
Project-Level Translation
In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program. This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design. This setting reflects how crochet instructions are used in practice, where each step depends on the correctness of all preceding steps. Image grounding is especially helpful for interpreting repeated motifs, symmetry, shaping, and termination conditions that may be under-specified in text alone.
@@
Because crochet patterns are inherently free-form (where multiple distinct programs can yield the same final product and a single natural-language instruction may admit several semantically equivalent DSL realizations), there is no canonical gold program for Task D. Exact string matching would therefore misjudge many correct solutions. Instead, CrochetBench evaluates correctness through functional executability using the CrochetPARADE validator, which checks whether a predicted DSL program is syntactically valid, structurally consistent, and fully executable.
We use two complementary evaluation settings. For step-level translation, we report the Valid Pattern Rate, defined as the proportion of generated DSL steps that successfully compile. For project-level translation, we compute the Valid Pattern Rate for full programs and, for those that compile, render the executable portion into a crochet-like image and compute its DINO Similarity (Oquab et al., 2023) to the ground-truth product image, providing a coarse measure of semantic fidelity beyond syntax. To diagnose failure modes, we further identify the first point of failure for each invalid prediction and categorize it using our fine-grained error taxonomy (Appendix E), enabling us to distinguish local symbolic errors from broader state-tracking failures or misinterpretations of the natural-language instruction.
Model Size Stitch Recognition (%) Instr. Sel. (%) Instr. Gen. (%)
Prec Rec F1 Acc BLEU R-L ChrF
Open Source BLIP-2 Flan-T5 XL 3B 29.53 23.03 22.50 25.62 0.22 9.30 9.45
Google Gemma 3 4B 20.54 10.21 12.65 24.94 0.11 3.67 5.59
Google Gemma 3 27B 17.19 18.14 16.05 24.94 0.40 5.17 6.55
DeepSeek-VL 7B 54.47 74.76 60.60 28.92 1.30 19.39 18.14
Qwen2-VL 7B 54.14 69.74 58.16 41.96 1.67 21.10 15.99
Qwen2-VL 72B 71.86 42.68 50.19 68.85 2.25 21.43 19.82
Closed Source GPT-4o 62.14 59.39 58.01 58.11 3.38 23.57 24.07
Gemini 2.5 Flash-Lite 74.49 49.77 56.83 55.63 4.93 25.92 30.50
Claude Sonnet 4 78.61 53.12 60.94 57.39 3.35 25.12 23.16
Finetuned Qwen2-VL (finetuned) 7B - - - - 5.64 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+3.97)$} 25.10 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+4.00)$} 22.39 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+6.40)$}
Table 2: Combined evaluation results across all three CrochetBench tasks: Stitch Recognition, Instruction Selection, and Instruction Generation. Best values are bold; second-best are underlined. For the finetuned Qwen2-VL model, values in parentheses indicate absolute improvements over the pretrained Qwen2-VL 7B baseline.
...A Midnight Blue ... ...B Copper Brown ... Ch4. Join with sl st to first ch to form ring ... 4th rnd: Sl st in next dc. Ch 2. 3 hdc in same dc as sl st. *Bobble in next st... Tassels (make 7)...
(a) Ground Truth
... With Color A (Blue), make a magic ring ... Join Color B (Brown) ... Rnd 4:Sl st into ch-2 sp. Ch 1, (Bobble ... Make 7 tassels with Color B and attach one to each ch-2 point of the star...
(b) Gemini
... Star-Shaped Crochet Blanket Pattern ... Yarn: Brown and Blue... With Brown yarn, make a magic ring. ... Tassels: Cut 4 lengths of Brown yarn ...
(c) GPT-4o
... 8-Point Star Afghan ... Worsted weight yarn in blue and brown ... With blue, ch 4, join with sl st to form ring. ... Tassels: Cut ...
(d) Claude
... Star Motif Crochet Pattern ... Worsted weight yarn in navy blue and brown ... Tassels... Round 1:Ch 10, join with sl st to form a ring ...
(e) Qwen-VL-7B
ch 100, sc in 2nd ch from hook, sc in each ch across, join with sl st in 1st sc, 100 loops made.
(f) DeepSeek VL
Figure 2: Case study for Task C: Instruction Generation. Each row shows the DSL-rendered output generated from the model’s natural-language instructions and the color-coded instruction extract below it. Matching colors denote semantically corresponding elements across the reference and model outputs, while red marks incorrect or hallucinated steps. The ground truth is a seven-point star with alternating blue and brown yarn and tassels attached at each point. Gemini and GPT-4o generate structured and mostly coherent instructions but misconstruct the global geometry, producing a circular motif rather than a star. Claude and Qwen2-VL-7B misinterpret the shape more severely, producing circular or distorted wireframe-like forms. DeepSeek-VL collapses entirely into a degenerate single-loop pattern. Gemini is the only model to explicitly recognize the motif as a seven-point star, but its instructions still fail to produce the correct star topology. To ensure transparency, we include the fully generated instructions for all models in Appendix F.
4 Experiments
We evaluate a representative set of widely used vision–language models spanning open and closed ecosystems. For open source models, we include BLIP-2 Flan-T5 XL (Li et al., 2023), Google Gemma 3 (4B and 27B) (Team et al., 2024), DeepSeek-VL 7B (Lu et al., 2024), and Qwen2-VL (7B and 72B) (Wang et al., 2024), covering a range of architectures and parameter scales. For closed source models, we evaluate GPT-4o (Hurst et al., 2024) , Gemini 2.5 Flash-Lite (Comanici et al., 2025), and Claude Sonnet 4 (Anthropic, 2025), which represent the strongest publicly accessible multimodal systems. These models span diverse architectures and parameter scales, providing a diverse and meaningful basis for assessing current multimodal capabilities on perception, retrieval, and procedural reasoning tasks.
We evaluate all models on the four CrochetBench tasks described in Section 3, using the same metrics and evaluation protocols throughout. The prompts used for evaluation are provided in Appendix C. Results and analysis are listed below.
Perception and grounding improve with scale, but procedural generation collapses.
Table 2 summarizes results across Stitch Recognition (Task A), Instruction Selection (Task B), and Instruction Generation (Task C). Closed-source models achieve the strongest recognition performance, with Claude Sonnet 4 obtaining the highest F1 score (60.94%), and DeepSeek-VL 7B leading among open models (60.60%). A clear precision–recall trade-off emerges across model scales: smaller models such as DeepSeek-VL 7B and Qwen2-VL 7B achieve higher recall but lower precision, whereas larger models such as Qwen2-VL 72B and closed-source models favor precision over recall. This pattern reflects different grounding strategies: smaller models tend to broadly enumerate common stitch types to maximize coverage, while larger models adopt a more selective, conservative approach, yielding higher precision but incomplete visual grounding.
Although larger models capture more fine-grained visual cues, accuracy remains far from saturated, and Instruction Selection shows similarly limited progress, with most large and closed-source models clustering around 55–60% accuracy, indicating that visual–textual alignment still depends on shallow correlations rather than robust grounding. These limitations become more pronounced in Task C. Natural-language instruction generation remains extremely challenging for all model, with BLEU, ROUGE-L, and ChrF scores uniformly low; even the strongest system, Gemini 2.5 Flash-Lite, achieves only 4.93% BLEU and 30.50% ChrF. The sharp drop from Tasks A and B to C shows that models capable of recognizing stitches or retrieving plausible text still fail to synthesize coherent multi-step procedures, reflecting fundamental gaps in procedural reasoning, symbolic consistency, and pattern-structure understanding.
Surface-level fluency does not imply procedural correctness.
Figure 2 presents a case study comparing model-generated natural-language instructions with their corresponding DSL renderings. Qwen2-VL-7B and DeepSeek-VL collapse into non-star geometries, revealing unstable procedural logic. GPT-4o and Claude produce structurally well-formed, human-like crochet instructions, including explicit materials sections and round-level organization, and correctly capture local yarn colors. However, they fundamentally misinterpret the global motif: GPT-4o reconstructs a four-point star and begins with the brown yarn instead of blue, while Claude generates an eight-point motif rather than the intended seven. Gemini most accurately identifies the seven-point structure and selects plausible constructs such as bobbles for the star tips, but structural inconsistencies remain and yield visibly distorted shapes. These examples demonstrate that models can generate fluent, crochet-like descriptions while failing to preserve the algorithmic structure required for faithful pattern synthesis.
Supervised finetuning improves fluency but not procedural correctness.
We evaluate supervised finetuning on CrochetBench-C using an 80/5/15 stratified train/validation/test split with full-parameter updates. As shown in Table 3 (bottom line), finetuning substantially improves all generation metrics: BLEU increases from 1.67% to 5.64% (+238%), ROUGE-L from 21.10% to 25.10% (+19%), and ChrF from 15.99% to 22.39% (+40%). The finetuned 7B model outperforms all evaluated closed-source systems, including GPT-4o, Claude Sonnet 4, and Gemini 2.5 Flash-Lite. Despite these gains, absolute performance remains low (5.64% BLEU), indicating that finetuning does not resolve core procedural failures. Qualitative analysis shows that improvements are largely superficial: outputs exhibit more consistent formatting, improved stitch vocabulary, and better adherence to crochet conventions, but still contain frequent stitch-count errors, malformed repeats, and incorrect global structure (e.g., rounds vs. rows, increases/decreases). These findings suggest that finetuning primarily enhances surface fluency rather than inducing robust algorithmic or structural reasoning.
Figure 3: Task D step-level translation results showing the proportion of generated DSL lines that successfully compile for early (Steps 1–2), middle (Steps 1–4), and late (Steps 1–6) stages of crochet patterns. Across all models, valid pattern rates increase as more context is provided, but overall accuracy remains low. Even the strongest models struggle in early steps, indicating difficulty establishing correct stitch state and structural dependencies. Larger models (e.g., Qwen2-VL-72B and Gemma-27B) do not consistently outperform their smaller counterparts, highlighting scale alone does not improve program-level structural reasoning.
Early-step instability reveals limits of procedural reasoning.
Figure 3 shows step-level results on Task D. Valid Pattern Rate increases with pattern depth but remains low overall: most models achieve under 15% validity in the first two steps, improve modestly in steps 3–4, and reach only 55–65% in later steps. This pattern reflects the difficulty of the initial steps, which must correctly initialize the program state such as defining stitch variables and maintaining balanced grouping. Errors made early propagate irreversibly, and later correctness often occurs only when the initial state is accidentally valid, indicating reliance on continuation heuristics rather than genuine procedural understanding. Larger models do not consistently perform better: Qwen2-VL-72B underperforms Qwen2-VL-7B, and Gemma-27B underperforms Gemma-4B, suggesting that increased capacity improves descriptive fluency more readily than symbolic stability, and that scaling alone is insufficient for grammar-sensitive procedural tasks.
Figure 4: Distribution of project-level DSL translation outcomes for each model, broken down into valid outputs and error categories. Across all models, invalid programs dominate, with most failures arising from undefined stitches, unbalanced brackets, and multiple-reference errors. The spread of error types further illustrates the difficulty of maintaining global consistency and symbolic correctness when generating full crochet programs.
Project-level DSL synthesis reveals severe long-range structural failures, and does not improve with model scale.
Figure 4 highlights the fragility of full-program CrochetPARADE synthesis. Valid outputs are rare: even the strongest models (Claude, Gemini, and Qwen2-VL-7B) produce only 5–8% executable programs, while most others fall below 3%. The most frequent failure modes—undefined stitches and unbalanced brackets—indicate poor control over the DSL vocabulary and grouping structure. Additional errors, including multiple references, non-adjacent labels, and runtime failures, further reveal difficulties in maintaining consistent state and long-range structural dependencies across an entire pattern.
Interestingly, scaling does not alleviate these issues and can exacerbate them. Larger models exhibit substantially higher rates of undefined-stitch errors, suggesting a tendency toward uncontrolled symbolic invention (e.g., Qwen2-VL-72B: 72.0% vs. Qwen2-VL-7B: 13.6%; Gemma-3-27B: 42.6% vs. Gemma-3-4B: 27.4%).
Image-based similarity confirms lack of global structural fidelity.
Compilation verifies syntactic and structural correctness but cannot determine whether two DSL programs are semantically equivalent. To address this gap, we compute DINO similarity between the target crochet product image and the rendering produced from each model’s executable program (valid outputs only). Figure 5 shows that similarity scores remain uniformly low across all models (0.10–0.17), far below the typical threshold for visually matched crochet images. Even when a model produces a compilable DSL program, the resulting rendering generally bears little resemblance to the intended pattern, indicating that syntactic validity does not imply correct procedural structure. The consistently low similarities reinforce that current multimodal LLMs fail to capture the global geometry and layout required for visually faithful crochet synthesis.
Figure 5: Task D Project-level translation results evaluating with DINO similarity between ground-truth images and DSL-rendered outputs generated from each model’s DSL program (valid executable portion only). The red line marks an approximate “good” similarity threshold. All models fall well below this level, indicating that even executable DSL programs rarely reproduce the correct visual structure of the intended crochet design.
5 Related Work
Multimodal benchmarks have traditionally emphasized descriptive image–text alignment, as in COCO (Lin et al., 2014) and Flickr30k (Plummer et al., 2015). More recent datasets extend to instructional or procedural understanding, including Recipe1M+ (Marin et al., 2018) and large-scale instructional video corpora such as YouCook2 and HowTo100M (Zhou et al., 2018; Miech et al., 2019). However, these benchmarks primarily evaluate semantic alignment, retrieval, or temporal recognition, and do not assess whether a model can generate or execute a correct procedure.
Texile crafts provide precisely such a domain. Crochet patterns specify symbolic, stepwise procedures that determine the topology and geometry of a final physical artifact. Prior work in this area including Digital Crochet (Seitz et al., 2022) and Neural Inverse Knitting (Kaspar et al., 2019) demonstrates the feasibility of representing textile structures in machine-readable form but remains limited in scale and modality. By moving from general instructional data to a structured craft domain, we enable multimodal models to be evaluated on artifact-centric procedural reasoning rather than temporal action recognition. CrochetBench builds on this emerging direction by providing thousands of real crochet patterns with paired images and natural-language instructions.
To evaluate procedural correctness, CrochetBench adopts an executable domain-specific language (CrochetPARADE), linking our tasks to program synthesis benchmarks such as HumanEval (Chen, 2021b), MBPP (Austin et al., 2021), and Spider (Yu et al., 2018). In multimodal settings, image-to-program benchmarks such as Im2LaTeX-100K (Deng et al., 2017) and pix2code (Beltramelli, 2018) similarly leverage executable formalisms for rendering-based evaluation. CrochetBench extends this executable perspective to textile crafts: patterns compile to structured instructions that can be rendered and validated, providing functional evaluation that tests whether a model’s output actually works. This offers a lightweight alternative to domains such as chemistry or cooking, where validating a procedure requires physical or chemical experiments that are slow, costly, or impractical to scale.
6 Conclusion
CrochetBench provides a structured benchmark for assessing whether multimodal LLMs can move from recognizing visual content to executing the step-by-step procedures required to produce a crochet pattern. Across all four tasks, models demonstrate a consistent gap: they can identify stitches and retrieve plausible instructions, but they fail to generate structurally valid procedures or produce executable programs. Qualitative analysis reveals that fluent outputs often misrepresent global structure, and scaling further exacerbates execution failures due to uncontrolled invention of stitch types. These results highlight the need for methods that target structured, long-horizon reasoning.
Limitations
While CrochetBench enables executable evaluation of visual procedural reasoning, it has several limitations. First, the benchmark focuses on a single creative domain, crochet, which although representative of structured symbolic procedures, may not capture all forms of procedural reasoning found in domains such as robotics, mechanical assembly, or scientific experimentation. As a result, generalization to other domains should be interpreted with caution.
Second, executable correctness is assessed through the CrochetPARADE domain specific language, which necessarily abstracts certain aspects of real world crochet practice. Multiple distinct procedures can produce visually similar artifacts, and some valid human written patterns may be penalized due to constraints of the language or limitations of the rendering process. Although we mitigate this issue using functional validation and image based similarity metrics, these signals remain imperfect proxies for true semantic equivalence.
Finally, while we evaluate a diverse set of state of the art vision language models, our finetuning experiments are limited to supervised learning on a single architecture. Exploring alternative training paradigms, such as program guided learning or learning with execution feedback, remains an important direction for future work.
Ethical Considerations